# Distributed computing quantum unitary evolution

### v2

@@ I Introduction
Complex quantum system modeling is one of the most important directions in computational mathematics today, especially in the computational fields involving polymer chemistry and macromolecular biology [1, 2, 3]. When simulating chemical and biological reactions, a large number of particles are often involved, and the dimension of the quantum system composed of these particles increase exponentially as the number of particles increases, thus causing a problem called the curse of dimensionality — raised by R.E. Bellman [4, 5] to describe a number of events that occur while organizing and evaluating data in high-dimensional areas. The curse of dimensionality has always been a major obstacle to the study of high-dimensional quantum systems. However, the study of the structure of biological macromolecules is one of the most important frontier studies of quantum computing, thus the curse of dimensionality is an urgent issue to be solved. With the development of supercomputers in recent decades, the use of distributed computing can solve a series of memory and efficiency problems caused by the curse of dimensionality to a certain extent. Moreover, some distributed computing algorithms can be applied to quantum unitary evolution. A key contribution of this paper is the cavity quantum electrodynamics (QED) model, which is easy to implement in the laboratory and offers a unique scientific paradigm for studying light-matter interaction. According to this paradigm, impurity two-level systems, also known as atoms, are connected to fields of cavities. The cavity QED model includes the Jaynes–Cummings model (JCM) [6] and the Tavis-Cummings model (TCM) [7] as well as their generalizations. Many studies have been conducted recently in the field of these models [8, 9, 10, 11, 12]. In our previous studies, some methods have been used to solve some of the computing problems caused by the curse of dimensionality [13, 14]. In this paper, we try to simulate the cavity QED model with a large number of two-level artificial atoms on a supercomputing platform to reveal the rules of unitary evolution of a complex quantum system, and study the time cost and speedup under different distributed computing strategies comparing with that under the situation without distributed computing. This paper is organized as follows. After introducing the TCM with a large number of artificial atoms in Sec. II, we introduce numerical method and its distributed computing transformation in Sec. III. We present the results of our numerical simulations in Sec. IV.1 and study the efficiency of distributed computing method in Sec. IV.2. Some brief comments on our results in Sec. V close out the paper.
@@ II Tavis–Cummings model with a large number of artificial atoms
$|\Psi\rangle=|p\rangle\bigotimes_{i=1}^{n}|l_{i}\rangle$ (1)
where $p\in[0,n]$ — number of free photons, $n$ — number of atoms, $l_{i},\ i\in[1,n]$ — electronic state, $l_{i}=0$ — i-st electron in ground state, $l_{i}=1$ — i-st electron in excited state. Since the closed system (dissipation process is disabled), thus we can omit the qubit involving photons
@@
And we can intuitively get that the dimension of Hilbert space is $N=2^{n}$. Interaction between atom and field is explained in detail in Fig. 1. In (a), an electron in the ground state absorbs a photon and transfer to the excited state, which is called excitation. In (b), an electron in the excited state transfer to the ground state after releasing a photon, this process is called de-excitation. Initial state is shown in Fig. 1 (c), where exist $n$ excited atoms and no photons. The excited atom will de-excite and become a ground state atom, at which time a photon will be released (see Fig. 1 (d)). When all atoms change to the ground state, $n$ photons are present in the optical cavity (see Fig. 1 (e)). In a closed system, the de-excitation and excitation of atom exist at the same time, so a certain regularity will appear in the light-matter system, which is embodied in the periodic oscillation of the quantum states. Before constructing the Hamiltonian, we first introduce rotating wave approximation (RWA) [15], which is taken into account
@@ III.2 Cannon’s algorithm
In this paper, we use Cannon’s algorithm [18] to solve the memory problem caused by large matrices with high dimension. This algorithm is used to process multiplication-addition operations of large matrices. Its core idea is to divide the matrix into many small blocks and send them to processors. Each processor receives its own block in sequence. This method can significantly reduce the occupied memory of each processor, but it also causes some disadvantages: as the number of processors increases, the time cost required for data transmission across processors and across nodes also increases. According to Eq. (2), the dimension of Hamiltonian and density matrix is equal to $N=2^{n}$. We suppose the number of processors is $p$, and $p_{x}=p_{y}=\sqrt{p}$. Thus, a matrix can be divided into $p_{x}\times p_{y}$ blocks. At the beginning, block on the processor $P_{i,j}$ is denoted by $A_{i,j}$ (or $B_{i,j}$), where $0\leq i,j<p_{x}$. The steps of Cannon’s algorithm are as follows
@@
If $M_{1}=M_{2}=\cdots=M_{k-1}=M_{k}=M$, $Cannon^{k}(M_{1},M_{2},\cdots,M_{k-1},M_{k})$ can be simply rewritten as $Cannon^{k}(M)$. After construction of Cannon’s algorithm, we can transform the Eq. (8) to a distributed computing form. In the first step, we get left matrix exponential part in Eq. (8) through Taylor series approximation
@@ IV.2 Comparison of cost time and speedup of different distributed computing strategies
We use the Cannon’s algorithm to solve distributed computing quantum unitary evolution on the supercomputer platform. We have four strategies of distributed computing: the first strategy is to split each matrix into $2\times 2=4$ blocks; the second strategy is to split each matrix into $4\times 4=16$ blocks; the third strategy — $8\times 8=64$ blocks; the forth strategy — $16\times 16=256$ blocks. And we compare the time cost and speedup of tasks under these strategies with the situation without distributed computing. The comparison of time cost of distributed computing Taylor series approximation is shown in Tab. 1 and the comparison of time cost of distributed computing unitary evolution is shown in Tab. 2. We found that when using $4$ computational cores for parallel computing, the time required to complete a task is always less than when no parallel computing strategy is used. However, the time cost of other strategies is more than that of the situation without distributed computing because of the high cost of data transmission between cores. However, as the dimension of the Hilbert space increases, the proportion of data transmission cost in the total time cost decreases, making the advantages of the strategy of high number of cores beginning to become obvious. Regarding the comparison of time cost, we can see it more intuitively in Fig. 5.

### v3

@@ I Introduction
Complex quantum system modeling is one of the most important directions in computational mathematics today, especially in the computational fields involving polymer chemistry and macromolecular biology [1, 2, 3]. When simulating chemical and biological reactions, a large number of particles are often involved, and the dimension of the quantum system composed of these particles increase exponentially as the number of particles increases, thus causing a problem called the curse of dimensionality — raised by R.E. Bellman [4, 5] to describe a number of events that occur while organizing and evaluating data in high-dimensional areas. The curse of dimensionality has always been a major obstacle to the study of high-dimensional quantum systems. However, the study of the structure of biological macromolecules is one of the most important frontier studies of quantum computing, thus the curse of dimensionality is an urgent issue to be solved. With the development of supercomputers in recent decades, the use of distributed computing can solve a series of memory and efficiency problems caused by the curse of dimensionality to a certain extent. Moreover, some distributed computing algorithms can be applied to quantum unitary evolution.
A key contribution of this paper is the cavity quantum electrodynamics (QED) model, which is easy to implement in the laboratory and offers a unique scientific paradigm for studying light-matter interaction. According to this paradigm, impurity two-level systems, also known as atoms, are connected to fields of cavities. The cavity QED model includes the Jaynes–Cummings model (JCM) [6] and the Tavis-Cummings model (TCM) [7] as well as their generalizations. Many studies have been conducted recently in the field of these models [8, 9, 10, 11, 12]. In our previous studies, some methods have been used to solve some of the computing problems caused by the curse of dimensionality [13, 14].
In this paper, we try to simulate the cavity QED model with a large number of two-level artificial atoms on a supercomputing platform to reveal the rules of unitary evolution of a complex quantum system, and study the time cost and speedup under different distributed computing strategies comparing with that under the situation without distributed computing. This paper is organized as follows. After introducing the TCM with a large number of artificial atoms in Sec. II, we introduce numerical method and its distributed computing transformation in Sec. III. We present the results of our numerical simulations in Sec. IV.1 and study the efficiency of distributed computing method in Sec. IV.2. Some brief comments on our results in Sec. V close out the paper.
@@ II Tavis–Cummings model with a large number of artificial atoms
$|\Psi\rangle=\bigotimes_{i=1}^{n}|p_{i}\rangle|l_{i}\rangle$ (1)
where $n$ — number of atoms, $p_{i}\in[0,1]$ — number of free photons from i-st atom, $l_{i}\in[0,1]$ — electronic state, $l_{i}=0$ — electron in ground state in i-st atom, $l_{i}=1$ — electron in excited state in i-st atom. Since the closed system, we can omit the qubit involving photons
@@
And we can intuitively get that the dimension of Hilbert space is $N=2^{n}$.
Interaction between atom and field is explained in detail in Fig. 1. In (a), an electron in the ground state absorbs a photon and transfer to the excited state, which is called excitation. In (b), an electron in the excited state transfer to the ground state after releasing a photon, this process is called de-excitation. Initial state is shown in Fig. 1 (c), where exist $n$ excited atoms and no photons. The excited atom will de-excite and become a ground state atom, at which time a photon will be released (see Fig. 1 (d)). When all atoms change to the ground state, $n$ photons are present in the optical cavity (see Fig. 1 (e)). In a closed system, the de-excitation and excitation of atom exist at the same time, so a certain regularity will appear in the light-matter system, which is embodied in the periodic oscillation of the quantum states.
Before constructing the Hamiltonian, we first introduce rotating wave approximation (RWA) [15], which is taken into account
@@ III.2 Cannon’s algorithm
In this paper, we use Cannon’s algorithm [18] to solve the memory problem caused by large matrices with high dimension. This algorithm is used to process multiplication-addition operations of large matrices. Its core idea is to divide the matrix into many small blocks and send them to processors. Each processor receives its own block in sequence. This method can significantly reduce the occupied memory of each processor, but it also causes some disadvantages: as the number of processors increases, the time cost required for data transmission across processors and across nodes also increases.
According to Eq. (2), the dimension of Hamiltonian and density matrix is equal to $N=2^{n}$. We suppose the number of processors is $p$, and $p_{x}=p_{y}=\sqrt{p}$. Thus, a matrix can be divided into $p_{x}\times p_{y}$ blocks. At the beginning, block on the processor $P_{i,j}$ is denoted by $A_{i,j}$ (or $B_{i,j}$), where $0\leq i,j<p_{x}$. The steps of Cannon’s algorithm are as follows
@@
If $M_{1}=M_{2}=\cdots=M_{k-1}=M_{k}=M$, $Cannon^{k}(M_{1},M_{2},\cdots,M_{k-1},M_{k})$ can be simply rewritten as $Cannon^{k}(M)$.
After construction of Cannon’s algorithm, we can transform the Eq. (8) to a distributed computing form. In the first step, we get left matrix exponential part in Eq. (8) through Taylor series approximation
@@ IV.1 The interaction between photons and matter
where $m$ is number of free photons in the cavity, $c_{i}$ — normalization factor. For example, $m=0$ means $\sum_{i=1}^{n}l_{i}=n-0=n$, that is to say, all atoms are in the excited state, that is, $l_{1}=l_{2}=\cdots=l_{n}=1$. If $m=n-1$, $\sum_{i=1}^{n}l_{i}=n-(n-1)=1$, that is to say, only one atom is in the excited state, and others are in the ground state.
@@ IV.2 Comparison of cost time and speedup of different distributed computing strategies
We use the Cannon’s algorithm to solve distributed computing quantum unitary evolution on the supercomputer platform. We have four strategies of distributed computing: the first strategy is to split each matrix into $2\times 2=4$ blocks; the second strategy is to split each matrix into $4\times 4=16$ blocks; the third strategy — $8\times 8=64$ blocks; the forth strategy — $16\times 16=256$ blocks. And we compare the time cost and speedup of tasks under these strategies with the situation without distributed computing.
The comparison of time cost of distributed computing Taylor series approximation is shown in Tab. 1 and the comparison of time cost of distributed computing unitary evolution is shown in Tab. 2. We found that when using $4$ computational cores for parallel computing, the time required to complete a task is always less than when no parallel computing strategy is used. However, the time cost of other strategies is more than that of the situation without distributed computing because of the high cost of data transmission between cores. However, as the dimension of the Hilbert space increases, the proportion of data transmission cost in the total time cost decreases, making the advantages of the strategy of high number of cores beginning to become obvious. Regarding the comparison of time cost, we can see it more intuitively in Fig. 5.